# Provable Scaling Laws of Feature Emergence from Learning Dynamics of Grokking

### v3

@@ Abstract
While the phenomenon of grokking, i.e., delayed generalization, has been studied extensively, it remains an open problem whether there is a mathematical framework that characterizes what kind of features will emerge, how and in which conditions it happens, and is still closely connected with the gradient dynamics of the training, for complex structured inputs. We propose a novel framework, named Li2, that captures three key stages for the grokking behavior of 2-layer nonlinear networks: (I) Lazy learning, (II) independent feature learning and (III) interactive feature learning. At the lazy learning stage, top layer overfits to random hidden representation and the model appears to memorize. Thanks to lazy learning and weight decay, the backpropagated gradient $G_{F}$ from the top layer now carries information about the target label, with a specific structure that enables each hidden node to learn their representation independently. Interestingly, the independent dynamics follows exactly the gradient ascent of an energy function $\mathcal{E}$, and its local maxima are precisely the emerging features. We study whether these local-optima induced features are generalizable, their representation power, and how they change on sample size, in group arithmetic tasks. When hidden nodes start to interact in the later stage of learning, we provably show how $G_{F}$ changes to focus on missing features that need to be learned. Our study sheds lights on roles played by key hyperparameters such as weight decay, learning rate and sample sizes in grokking, leads to provable scaling laws of feature emergence, memorization and generalization, and reveals the underlying cause why recent optimizers such as Muon can be effective, from the first principles of gradient dynamics. Our analysis can be extended to multi-layer architectures.
@@ 1 Introduction
Modeling the memorization/generalization behaviors have been a goal of many works. One such behavior, know as grokking (Power et al., 2022; Doshi et al., 2024; Nanda et al., 2023; Wang et al., 2024a; Varma et al., 2023; Liu et al., 2023; Thilak et al., 2022), shows that the model initially overfits to the training set, and then suddenly generalizes to unseen test samples after continuous training. Many explanation exists, e.g., effective theory (Liu et al., 2022; Clauw et al., 2024), efficiency of memorization and generalization circuits (Varma et al., 2023), Bayesian interpretation with weight decay as prior (Millidge, 2022), etc. Most works focus on a direct explanation of its empirical behaviors, or leveraging property of very wide networks (Barak et al., 2022; Mohamadi et al., 2024; Rubin et al., 2024), but few explores the details of the grokking learning procedure by studying the gradient dynamics on the weights.
@@
  • Explanation of Grokking behavior. At the beginning of grokking, the lazy learning corresponds to memorization, in which the top layer finds a (temporary) solution to explain the target with the random features. Only after that, the backpropagated gradient $G_{F}$ becomes meaningful, and causes the hidden layer to learn generalizable, emerging features.
@@ 2 Problem formulation
Previous works pointed out that grokking mostly happens when there is regularization during training (e.g., weight decay (Power et al., 2022; Nanda et al., 2023), Jacobian regularization (Walker et al., 2025), etc.). It remains a mystery why this is the case. In this work, we show that grokking is a consequence of “leaked” backpropagated gradient due to regularization.
Refer to captionRefer to caption
@@ 3 Stage I: Lazy Learning (Overfitting)
where $P_{\eta}=I-\tilde{F}(\tilde{F}^{\top}\tilde{F}+\eta I)^{-1}\tilde{F}^{\top}=\eta(\tilde{F}\tilde{F}^{\top}+\eta I)^{-1}$ (by Woodbury matrix formula). Note that $P_{\eta}\tilde{Y}=\tilde{Y}-\tilde{F}V_{\textrm{ridge}}$, which is the error of the output layer. Without weight decay (i.e., $\eta=0$), if the network is overparameterized and we have enough random features, then $P_{\eta}\tilde{Y}=\tilde{Y}-\tilde{F}V_{\textrm{ridge}}=0$ and thus $G_{F}=0$. In this case, feature learning does not happen (the bottom row of Fig. 3). Note that this does not rule out the possibility that feature learning happens during the period that $V$ converges to $V_{\textrm{ridge}}$, even if $\eta=0$ (Kumar et al., 2024). This is possible in particular if the (hidden) weights are initialized large (Clauw et al., 2024).
@@ 4.1 The energy function $\mathcal{E}$
Assume that (1) entries of $W$ follow normal distribution, (2) $\|\mathbf{x}_{i}\|_{2}$ have the same norm, (3) $\mathbf{x}_{i}^{\top}\mathbf{x}_{i^{\prime}}=\rho$ for all $i\neq i^{\prime}$ and (4) large width $K$, then both $\tilde{F}^{\top}\tilde{F}$ and $\tilde{F}\tilde{F}^{\top}$ becomes a multiple of identity and Eqn. 5 becomes:
$G_{F}=\frac{\eta}{(Kc_{1}+\eta)(nc_{2}+\eta)}\tilde{Y}\tilde{Y}^{\top}F$ (6)
where $c_{1},c_{2}>0$ are constants related to nonlinearity. When $\eta$ is small, we have $G_{F}\propto\eta\tilde{Y}\tilde{Y}^{\top}F$.
@@ 4.4 Representation power of learned features
Assumptions of the theorem. Training the model in an end-to-end manner automatically satisfies the assumption (3). If we initialize the weights randomly, then with high probability, the resulting $\mathbf{u}$ are not collinear and the assumption (2) can be satisfied. For assumption (1), interestingly, no such theorem can be stated in the real domain $\mathbb{R}$, because in $\mathbb{R}$, the subspace of orthogonal matrices, which group representations belong to, is not covered by the subspace of symmetric matrices spanned by $\{\mathbf{u}\mathbf{u}^{\top}\}$. In contrast, in the complex domain $\mathbb{C}$, the subspace of unitary matrices can be represented by Hermitian matrices. Does that mean that the real representation is not that useful? Not really. If we change $\mathbf{w}=[\mathbf{u};\pm P\mathbf{u}]$ slightly to $\mathbf{w}=[\mathbf{u};\pm P\mathbf{u}^{\prime}]$ in which $\mathbf{u}^{\prime}$ is a small perturbation of $\mathbf{u}$, then Thm. 3 holds for real solutions. This happens in the stage III when end-to-end backpropagation refines the representation.
@@ 9 Conclusion
If $s_{i}\equiv 1$ and $\rho_{ij}\equiv\rho$ for $i\neq j$, then
$\mathcal{K}(i,i)=\sum_{l\geq 0}a_{l}(s)^{2}=:a,\qquad\mathcal{K}(i,j)=\sum_{l\geq 0}a_{l}(s)^{2}\,\rho^{\,l}=:b,\quad(i\neq j),$
hence $\mathcal{K}(i,j)=(a-b)\delta_{ij}+b$ and thus $FF^{\top}=K(a-b)I+Kb\mathbf{1}\mathbf{1}^{\top}$. Note that by Parseval’s identity, $a=\mathbb{E}_{Z\sim\mathcal{N}(0,1)}[\sigma^{2}(sZ)]$.
Therefore, $\tilde{F}\tilde{F}^{\top}=K(a-b)P^{\perp}_{1}=K(a-b)(I-\mathbf{1}\mathbf{1}^{\top}/n)$ and $P_{\eta}\tilde{Y}=\frac{\eta}{K(a-b)+\eta}\tilde{Y}$. Since $\tilde{F}^{\top}\tilde{F}$ is proportional to identity matrix, $(\tilde{F}^{\top}\tilde{F}+\eta I)^{-1}$ is also proportional to identity matrix and the conclusion follows. ∎

### v4

@@ Abstract
While the phenomenon of grokking, i.e., delayed generalization, has been studied extensively, it remains an open problem whether there is a mathematical framework that characterizes what kind of features will emerge, how and in which conditions it happens, and is still closely connected with the gradient dynamics of the training, for complex structured inputs. We propose a novel framework, named Li2, that captures three key stages for the grokking behavior of 2-layer nonlinear networks: (I) Lazy learning, (II) independent feature learning and (III) interactive feature learning. At the lazy learning stage, top layer overfits to random hidden representation and the model appears to memorize. Thanks to lazy learning and weight decay, the backpropagated gradient $G_{F}$ from the top layer now carries information about the target label, with a specific structure that enables each hidden node to learn their representation independently. Interestingly, the independent dynamics follows exactly the gradient ascent of an energy function $\mathcal{E}$, and its local maxima are precisely the emerging features. We study whether these local-optima induced features are generalizable, their representation power, and how they change on sample size, in group arithmetic tasks. When hidden nodes start to interact in the later stage of learning, we provably show how $G_{F}$ changes to focus on missing features that need to be learned. Our study sheds lights on roles played by key hyperparameters such as weight decay, learning rate and sample sizes in grokking, leads to provable scaling laws of feature emergence, memorization and generalization, and reveals the underlying cause why recent optimizers such as Muon can be effective, from the first principles of gradient dynamics. Our analysis can be extended to multi-layer architectures. The code is available [^1] https://github.com/yuandong-tian/understanding/tree/main/ssl/real-dataset/cogo.
@@ 1 Introduction
Modeling the memorization and generalization behaviors have been a goal of many works. One such behavior, know as grokking (Power et al., 2022; Doshi et al., 2024; Nanda et al., 2023; Wang et al., 2024a; Varma et al., 2023; Liu et al., 2023; Thilak et al., 2022), shows that the model initially overfits to the training set, and then suddenly generalizes to unseen test samples after continuous training. Many explanation exists, e.g., effective theory (Liu et al., 2022; Clauw et al., 2024), efficiency of memorization and generalization circuits (Varma et al., 2023), Bayesian interpretation with weight decay as prior (Millidge, 2022), etc. Most works focus on a direct explanation of its empirical behaviors, or leveraging property of very wide networks (Barak et al., 2022; Mohamadi et al., 2024; Rubin et al., 2024), but few explores the details of the grokking learning procedure by studying the gradient dynamics on the weights.
@@
  • Explanation of Grokking behavior. At the beginning of grokking, the lazy learning corresponds to memorization, in which the top layer finds a (temporary) solution to explain the target with the random features. Only after that, the backpropagated gradient $G_{F}$ becomes meaningful, and causes the hidden layer to learn generalizable, emerging features. Grokking is accelerated with regularization.
@@ 2 Problem formulation
While grokking can happen without regularization (Gromov, 2023), it is often stabilized and accelerated with it (e.g., weight decay (Power et al., 2022; Nanda et al., 2023), Jacobian regularization (Walker et al., 2025), etc.). In our framework, we incorporate weight decay and show that grokking is a consequence of “leaked” backpropagated gradient due to regularization.
@@ 3 Stage I: Lazy Learning (Overfitting)
where $P_{\eta}=I-\tilde{F}(\tilde{F}^{\top}\tilde{F}+\eta I)^{-1}\tilde{F}^{\top}=\eta(\tilde{F}\tilde{F}^{\top}+\eta I)^{-1}$ (by Woodbury matrix formula). Note that $P_{\eta}\tilde{Y}=\tilde{Y}-\tilde{F}V_{\textrm{ridge}}$, which is the error of the output layer. Without weight decay (i.e., $\eta=0$), if the network is overparameterized and we have enough random features, then $P_{\eta}\tilde{Y}=\tilde{Y}-\tilde{F}V_{\textrm{ridge}}=0$ and thus $G_{F}=0$. In this case, feature learning does not happen (the bottom row of Fig. 3). Note that this does not rule out the possibility that feature learning may happen during the period that $V$ converges to $V_{\textrm{ridge}}$, even if $\eta=0$ (Kumar et al., 2024; Gromov, 2023). This is possible in particular if the (hidden) weights are initialized large (Clauw et al., 2024).
Fig. 12 and Fig. 13 shows that using $V_{\mathrm{ridge}}$ during training (rather than using gradient descent to update $V$) still leads to grokking and feature learning. Moreover, it groks faster for larger $M$. This shows that replacing gradient update of $V$ with $V_{\mathrm{ridge}}$ may not change the feature learning dynamics, and at the same time makes it easier to analyze.
@@ 4.1 The energy function $\mathcal{E}$
Assume that (1) entries of $W$ follow standard normal distribution $N(0,1)$, (2) $\|\mathbf{x}_{i}\|_{2}=\text{const}$, (3) $\|\mathbf{x}_{i}^{\top}\mathbf{x}_{i^{\prime}}-\rho\|_{2}\leq\epsilon$ for all $i\neq i^{\prime}$ and (4) large width $K$, then both $\tilde{F}^{\top}\tilde{F}$ and $\tilde{F}\tilde{F}^{\top}$ becomes a multiple of identity and Eqn. 5 becomes:
$G_{F}=\frac{\eta}{(Kc_{1}+\eta)(nc_{2}+\eta)}\tilde{Y}\tilde{Y}^{\top}F+O(K^{-1}\epsilon)$ (6)
where $c_{1},c_{2}>0$ are constants related to nonlinearity. When $\eta$ is small, we have $G_{F}\propto\eta\tilde{Y}\tilde{Y}^{\top}F$. Note that the input features and/or weights can be scaled and what changes is $c_{1}$ and $c_{2}$.
@@ 4.4 Representation power of learned features
Assumptions of the theorem. Training the model in an end-to-end manner automatically satisfies the assumption (3). If we initialize the weights randomly, then with high probability, the resulting $\mathbf{u}$ are not collinear and the assumption (2) can be satisfied. For assumption (1), interestingly, no such theorem can be stated in the real domain $\mathbb{R}$, because in $\mathbb{R}$, the subspace of orthogonal matrices, which group representations belong to, is not covered by the subspace of symmetric matrices spanned by $\{\mathbf{u}\mathbf{u}^{\top}\}$. In contrast, in the complex domain $\mathbb{C}$, the subspace of unitary matrices can be represented by Hermitian matrices. Does that mean that the real representation is not that useful? Not really. If we change $\mathbf{w}=[\mathbf{u};\pm P\mathbf{u}]$ slightly to $\mathbf{w}=[\mathbf{u};\pm P\mathbf{u}^{\prime}]$ in which $\mathbf{u}^{\prime}$ is a small perturbation of $\mathbf{u}$, then Thm. 3 holds for real solutions. This happens in the stage III when end-to-end backpropagation refines the representation. Fig. 11 shows that using complex weights still works and may grok faster.
@@ 7 Discussion
Relationship between grokking and regularization. A key insight of the Li2 framework is that a non-zero backpropagated gradient $G_{F}$ initiates the feature learning process, setting the stage for optimizing the energy function $\mathcal{E}$ that determines which features are learned and how. Intuitively, as long as the lazy learning phase produces a sufficiently large nonzero $G_{F}$ correlated with the target label $Y$, feature learning will be triggered.
Several ways exist to ensure this nonzero gradient. Regularization mechanisms, such as weight decay, help guarantee the presence of a nonzero $G_{F}$, and conveniently allow us to demonstrate the feature learning process in a mathematically rigorous way. However, we want to emphasize that regularization is not the only option. For example, $G_{F}$ might become large during lazy learning and later vanish due to perfect overfitting (see Eqn. 5). Notably, (Prieto et al., 2024) replaces the standard softmax (exponential term) with a stable softmax (linear term), which from the perspective of the Li2 framework, slows down overfitting to the labels. This maintains a nonzero backpropagated gradient, thereby preserving the potential for subsequent feature learning.
In summary, our framework shows that weight decay can be a sufficient condition for feature learning (and grokking), but not a necessary one.
@@ 9 Conclusion, Limitations, and Future Work
Limitations. While the derivation of energy $\mathcal{E}$ is applicable to any input, analysis of its local maxima relies on restrictive assumption of group structure of the input. Also our analysis does not include the transition time between consecutive learning stages. We leave them for future work.
@@
If $s_{i}\equiv s$ and $\|\rho_{ij}-\rho\|_{2}\leq\epsilon$ for $i\neq j$, then
$\mathcal{K}(i,i)=\sum_{l\geq 0}a^{2}_{l}(s)=:a$
Let $c:=\sum_{l\geq 1}la^{2}_{l}(s)<+\infty$ (it is convergent due to the big factor $l!$ in the denominator). Let $b:=\sum_{l\geq 0}a^{2}_{l}(s)\,\rho^{\,l}$ and we have for all $i\neq j$:
$\|\mathcal{K}(i,j)-b\|_{2}\leq\sum_{l\geq 0}a^{2}_{l}(s)\,\|\rho^{\,l}_{ij}-\rho^{l}\|_{2}\leq\sum_{l\geq 1}la^{2}_{l}(s)\,\epsilon=c\epsilon$
due to the fact that $\|\rho^{\,l}_{ij}-\rho^{l}\|_{2}\leq l\xi^{l-1}\epsilon$ for all $l\geq 1$ and some $\xi$ in between $\rho_{ij}$ and $\rho$. hence $\mathcal{K}(i,j)=(a-b)\delta_{ij}+b+O(\epsilon)$ and thus $FF^{\top}=K(a-b)I+Kb\mathbf{1}\mathbf{1}^{\top}+O(K\epsilon)\mathbf{1}\mathbf{1}^{\top}$. Note that by Parseval’s identity, $a=\mathbb{E}_{Z\sim\mathcal{N}(0,1)}[\sigma^{2}(sZ)]$.
Therefore, $\tilde{F}\tilde{F}^{\top}=K(a-b+O(\epsilon))P^{\perp}_{1}=K(a-b+O(\epsilon))(I-\mathbf{1}\mathbf{1}^{\top}/n)+O(K\epsilon)\mathbf{1}\mathbf{1}^{\top}$ and $P_{\eta}\tilde{Y}=\frac{\eta}{K(a-b)+\eta}\tilde{Y}$. Since $\tilde{F}^{\top}\tilde{F}$ is proportional to identity matrix, $(\tilde{F}^{\top}\tilde{F}+\eta I)^{-1}$ is also proportional to identity matrix and the conclusion follows. ∎
@@ C.1 Use Groups Algorithms Programming (GAP) to get non-Abelian groups
Figure 10: Training modular addition tasks with real weights ($M=23,41,89,127$). Learning rate is $0.005$, weight decay is 5e55e-5. Number of hidden nodes $K=256$. Test sample is $20\%$ of the full set of $M^{2}$. Using Adam optimizer. Averaged over 55 seeds. This is a baseline.
Figure 11: Training modular addition tasks with complex weights ($M=23,41,89,127$). Learning rate is $0.005$, weight decay is 5e55e-5. Number of hidden nodes $K=256$. Test sample is $20\%$ of the full set of $M^{2}$. Using Adam optimizer. Averaged over 55 seeds. Compared with the real case (Fig. 10), models with complex weights seem to grok faster.
Figure 12: Training modular addition tasks with real weights ($M=23,41,89,127$). Instead of using gradient descent to update the top layer $V$, in every gradient update we use ridge regression solution $V_{\textrm{ridge}}$ with respect to the current $F$ (Eqn. 4). Learning rate is $0.005$, weight decay is 5e55e-5. Number of hidden nodes $K=256$. Test sample is $20\%$ of the full set of $M^{2}$. Using Adam optimizer. Averaged over 55 seeds. The grokking still happens (for $M=23$ check Fig. 13 for completeness). It is slower for $M=23$ but actually faster for $M=41,89,127$, compared to the baseline (Fig. 10).
Figure 13: Training modular addition tasks with real weights $M=23$ for 500 epochs, using $V_{\mathrm{ridge}}$ as the top layer weight. The grokking still happens but slower than the baseline (Fig. 10) for $M=23$.